I do not know what kind of problem this is but i need help.

the sum of three of three consectutive even integers is 174. Find the integers

Answer by stanbon(75887)

the sum of three consectutive even integers is 174. Find the integers.
--------------------------
1st: 2x
2nd: 2x+2
3rd: 2x+4
--------------
Equation:
2x + 2x+2 + 2x+4 = 174
6x+6 = 174
6x = 168
x = 28
---
1st: 2*28 = 56
2nd: 58
3rd: 60
==================
Cheers,
Stan H.